# Twin-field-based multi-party quantum key agreement

### v1

@@ Abstract
Quantum key distribution (QKD) can secure cryptographic communication between two distant users, as guaranteed by the laws of quantum mechanics rather than computational assumptions. The twin-field scheme, which employs counter-propagated weak coherent light pulses, doubles the secure distance of standard QKD without using quantum repeaters. Here, we study a method to extend the twin-field key distribution protocol to a scheme for multi-party quantum key agreement. We study our protocol’s security using a minimum error discrimination analysis and derive the asymptotic key rate based on the entanglement-based source-replacement scheme. We also simulate it on the ANSYS Interconnect platform to study the protocol’s performance in certain practical situations.
@@ 1 Introduction
The ability of quantum key distribution (QKD) [1, 2] to provide two remote users, Alice and Bob, with a secret key that is both composable and unconditionally secure with the principles of quantum mechanics is remarkable in quantum communication and cryptography [3, 4, 5, 6]. Recently, there has been development in practical implementations of secure quantum communication across longer distances due to novel twin-field QKD protocol [7]. The maximum key rate for traditional point-to-point QKD systems increases linearly with the channel transmittance $\eta$ based on the losses [8, 9]. Twin-field QKD offers a better rate-loss tradeoff since the key rate grows linearly with $\sqrt{\eta}$ [10].
@@
A twin-field QKD network would have a remarkable advantage over existing networks in terms of secure key rate, and it is intrinsically a measurement-device-independent QKD impervious to side-channel attacks [20]. A Sagnac interferometer-based twin-field QKD for multiple parties was implemented recently for pairwise communication [21, 22, 23]. The twin-field scheme can also be applied to standard quantum cryptography protocols employing single photons or entangled states, for example, quantum digital signature [24] constructed on the basis of the counterfactual principle [25, 26].
In this spirit, we propose a technique for carrying out a multi-party quantum key agreement procedure in which three parties share a key concurrently using the twin-field protocol. The remaining paper is sectioned as follows. The protocol is described in Sec. 2.1 with minimum error discrimination analysis against individual attacks in Sec. 2.2. The asymptotic key rate based on security against collective attacks is demonstrated in Sec. 3. Using simulations is a useful way to comprehend the workings of the QKD systems and the protocol implemented. The simulations for the proposed multi-party quantum key agreement procedure are performed on the ANSYS Interconnect platform, and the details are given in Sec. 3.1. Some difficulties in the practical implementation of this protocol are discussed in Sec. 4.
@@ 2 Multi-party quantum key agreement based on twin-field cryptography
2.1 Description of the protocol
In the multi-party twin-field protocol, Alice, Bob, and Charlie send information to two nodes, AB and BC, as shown in Fig. 1. For $N$-party protocol, $N-1$ nodes are required as per this conception.
@@ 2.1 Description of the protocol
      At BC node, if $\delta_{2}=+$, Bob and Charlie store their respective outcomes as the shared bits. If $\delta_{2}=-$, Charlie flips his bit value. Repeating this, Charlie and Bob obtain key $k_{BC}$.
@@
      Suppose $|k_{AB}|=|k_{BC}|$. Bob announces $k_{AB}\oplus k_{BC}$, using which using which Alice and Charlie individually retrieve $k_{BC}$ and $k_{AB}$. If $|k_{AB}|\neq|k_{BC}|$, then $||k_{AB}|-|k_{BC}||$ bits are randomly removed from the longer key, and the above procedure is applied.
@@
2.2 Minimum error discrimination on the multi-party twin-field protocol
Initially we perform a simple minimum error discrimination analysis (on individual pulses) on the multi-party protocol to show its robustness compared to two-party protocol. Assuming that Eve controls the nodes AB, BC and the channels $l_{A}$, $l_{B}$, $l^{\prime}_{B}$ and $l_{C}$, she makes the announcements $\delta_{1}(\delta_{2})$ [27, 28]. Eve must effectively discriminate between the joint states that are correlated vs anti-correlated at node AB
@@ 2.2 Minimum error discrimination on the multi-party twin-field protocol
3 Security and key rate for the multi-party twin-field protocol
To demonstrate the security and asymptotic key rate of the twin-field protocol, Ref. [28] introduces the source-replacement strategy [30], which is applied to Alice and Bob’s sources. Here we extend this to Charlie’s source. Then, we adapt the multi-party protocol to its equivalent entanglement-based protocol to verify the security against collective attacks in the asymptotic-key scenario. The key rate is then assessed to demonstrate the entanglement-based version’s security.
@@ 3 Security and key rate for the multi-party twin-field protocol
3.1 Simulations of the multi-party twin-field protocol for key rate
@@
We perform simulations of the protocol for key rate analysis and describe the practical implementation. The image in Fig. 2 shows a simulation of the ANSYS Interconnect platform for the multi-party twin-field protocol defined in Sec. 2.1. The rectangular selections show the modules of Alice, Bob and Charlie from top to bottom on the left side of Fig. 2.
@@
We analyse the key rate by performing multiple iteration of the simulations for varied distance $L$ with losses described by Eq. (29) and obtain the key rate which closely matches with the result from previous section.
@@ 4 Difficulty in practical implementation
Practically putting in place a twin-field protocol between multiple parties presents technological difficulties since active phase stabilisation is necessary to interfere with coherent pulses due to random phase fluctuations [10, 34, 35, 36]. The central nodes AB and BC function as gigantic Mach-Zehnder interferometers for interference measurement, as shown in Fig. 1. Stabilising an unbalanced Mach-Zehnder interferometer with widely different path lengths is challenging. So, the setup would necessitate that users Alice-Bob (Bob-Charlie) have an identical geographic distance to the central node AB(BC) for the interference of pulses ideally, whereas, in the real world, there could be significant variations in the geographic distances between various parties and the nodes. A more practical method to accomplish this is via sending-or-not-sending, a unique type of twin-field protocol where data is encoded based on the decision to send in a coherent state (sending) or a vacuum state (not sending) and it has tolerance to phase misalignment errors [37].
@@ 5 Conclusion
We propose and examine a technique that expands on the two-party twin-field quantum key distribution to obtain a protocol for the concurrent multi-party quantum key agreement. For the proposed protocol, we carry out the asymptotic key rate analysis from a security proof based on the technique of entanglement-based source replacement. We simulated the protocol on the ANSYS Interconnect platform by assessing its practical implementation. These simulations can be augmented by including realistic experimental issues such as phase misalignment errors and the use of decoy states with varying intensities. This exercise can be extended to extensively study the variant of the present protocol based on the sending-or-not-sending twin-field protocol [37]. These simulations can also be used to quantify the information lost by the quantum hacking efforts [3, 5, 39] of an eavesdropper by exploiting device imperfections [40].

### v2

@@ Abstract
Quantum key distribution (QKD) can secure cryptographic communication between two distant users, as guaranteed by the laws of quantum mechanics rather than computational assumptions. The twin-field scheme, which employs counter-propagated weak coherent light pulses, doubles the secure distance of standard QKD without using quantum repeaters. Here, we study a method to extend the twin-field key distribution protocol to a scheme for multi-party quantum key agreement. We study our protocol’s security using a minimum error discrimination analysis and derive the asymptotic key rate based on the entanglement-based source-replacement scheme. We also simulate it on the ANSYS Interconnect platform with optical components to study the protocol’s performance in certain practical situations.
@@ 1 Introduction
The ability of quantum key distribution (QKD) [1, 2] to provide two remote users, Alice and Bob, with a secret key that is both composable and unconditionally secure with the principles of quantum mechanics is remarkable in quantum communication and cryptography [3, 4, 5, 6]. Recently, there has been development in practical implementations of secure quantum communication across longer distances due to novel twin-field QKD protocol [7]. The maximum key rate for traditional point-to-point QKD systems increases linearly with the channel transmittance $\eta$ based on the losses [8, 9]. Twin-field QKD offers a better rate-loss tradeoff since the key rate grows linearly with $\sqrt{\eta}$ [10, 11, 12].
@@
A twin-field QKD network would have a remarkable advantage over existing networks in terms of secure key rate, and it is intrinsically a measurement-device-independent QKD impervious to some side-channel attacks [22, 23]. A Sagnac interferometer-based twin-field QKD for multiple parties was implemented recently for pairwise communication [24, 25, 26]. The twin-field scheme can also be applied to standard quantum cryptography protocols employing single photons or entangled states, for example, quantum digital signature [27] constructed on the basis of the counterfactual principle [28, 29].
Quantum key agreement protocols attempt to share secret keys among several parties [30, 31, 32, 33]. It has been demonstrated that quantum key agreement has advantages compared to repeated QKD in quantum networks [34]. Many quantum key agreement protocols have been suggested [35], including measurement-device-independent [36] and device-independent [37] versions. However, similar to QKD, $\eta$ restricts the key rates in most quantum key agreement protocols [38]. Twin-field QKD has been implemented in quantum key agreement protocols for symmetric [39, 40, 41, 42] and asymmetric quantum channels [43] to break this constraint.
In this spirit, we propose a simple technique for a multi-party quantum key agreement procedure in which three parties share a key concurrently using the twin-field protocol, which can be straightforwardly extended for more parties. Extending twin-field to multi-field is not feasible as the experimental complications such as phase stabilization get compounded [10, 44, 45, 46]. Thus, we reduce the problem to a twin-field implementation where existing technology can be used for a higher number of parties. Moreover, this technique works for the arbitrary topology of a practical network, not just cyclic Sagnac interferometer-based twin-field [24, 25, 26].
The paper is sectioned as follows: The protocol is described in Sec. 2 with minimum error discrimination analysis against individual attacks in Sec. 3. The asymptotic key rate based on security against collective attacks is demonstrated in Sec. 4. The simulations with optical components for the proposed multi-party quantum key agreement procedure are performed on the ANSYS Interconnect platform and the details are given in Sec. 5. Further, some difficulties in the practical implementation of this protocol are discussed in Sec. 6.
@@ 2 Multi-party quantum key agreement based on twin-field cryptography
In multiparty quantum key agreement protocols the key rate for $N$ parties generally suffers as O($\eta^{N}$) or O($\eta^{N-1}$) [40] and extending the twin-field implementation for $N>3$ parties is non-trivial [41, 42, 43]. Current technology does not facilitate the experimental demonstration of the scheme implemented in Ref. [39]. Even considering a triple-field interference (where Alice, Bob, and Charlie send pulses to a central hub) such as in Fig. 1 would create large misalignment issues, and because the combinatorics of participants’ choice grows exponentially with N, it would generally not be favorable.
Figure 1: Possible tri-party triple-field implementation of quantum key agreement: three-mode linear interferometer injected with three weak coherent pulses.
Thus, we are led to consider the problem of extending the twin-field scheme to the multiparty case so that the standard twin-field protocol remains a fixed subroutine applied by different sets of parties spread out over a wide area. Our simple but fundamental observation here is that key agreement (as against key distribution) may circumvent the distance factor because it involves multiple trusted parties, making it likely that any given party will always find another involved in the close neighborhood. Our protocol exploits this property to optimize the quantum resources used. We begin by considering a three-party protocol.
2.1 Description of the three-party protocol
In the three-party twin-field protocol, Alice, Bob, and Charlie send information to two nodes, AB and BC, as shown in Fig. 2. For $N$-party protocol, $N-1$ nodes are required as per this conception.
@@ 2.1 Description of the three-party protocol
      If $\delta_{2}=+$ at the BC node, Bob and Charlie store their respective outcomes as the shared bits. If $\delta_{2}=-$, Charlie flips his bit value. Repeating this, Charlie and Bob obtain key $k_{BC}$.
@@
      Suppose $|k_{AB}|=|k_{BC}|$. Bob publicly announces $k_{AB+BC}\equiv k_{AB}\oplus k_{BC}$. Using her (resp., his) private copy of $k_{AB}$ (resp., $k_{BC})$, Alice (resp., Charlie) retrieves $k_{AB}=k_{BC}\oplus k_{AB+BC}$ (resp., $k_{BC}=k_{AB}\oplus k_{AB+BC}$). If $|k_{AB}|\neq|k_{BC}|$, then $||k_{AB}|-|k_{BC}||$ bits are removed by public discussion from the longer key, and the above procedure is applied.
@@
To generalize the above protocol, suppose we are given a group of $N>3$ parties that must agree on a key. First, a classical optimization algorithm is run to determine the optimal network that links all parties while minimizing the total length of all edges. Barring material or cost obstacles (which would be additional constraints, if applicable), this optimal network will constitute the required quantum network. The network’s vertices are divided into $n$ segments of three connected vertices, so each segment shares a vertex with at least one other segment. Ideally, we have $2n+1=\lceil N\rceil$, where the ceiling function represents round off to the following highest odd number (cf. Fig. 3).
Figure 3: Network with $N=7$ parties and $n=3$ twin-field key agreement implementations, among the parties $\{1,3,2\},\{3,4,5\}$ and $\{6,5,7\}$. Public announcements are made at vertices 3, 4 and 5. The key rate does not scale linearly with the size of the network (identified with the maximum distance between two parties, in this case, $d_{1,7}$), but instead quadratically between the longest nearest-neighbor distance, in this case, $d_{2,3}$.
Each segment will run the twin-field quantum key agreement protocol described above. The keys obtained in individual segments can be reconciled by inter-segmental public announcements by the parties shared between segments (different from but analogous to the intra-segmental announcements mentioned above). Suppose the keys (assumed to be of equal size) obtained from two contiguous twin-field segments are $s_{1}$ (for instance, $K$ in the above description) and $s_{2}$. Then, the XOR of the two keys, $s_{12}\equiv s_{1}\oplus s_{2}$, is announced. The remaining members of the first segment can derive $s_{2}=s_{12}\oplus s_{1}$, similarly for the other segment. The number $m$ of inter-segmental announcing parties generally depends on the network topology. For the circuit in Fig. 3, $m=3$, whereas for a network with star topology (a hub vertex, with the rest radiating from it), $m=1$ corresponds to the announcing party at the hub.
A practical consequence of our approach is that the key rate in our protocol will scale as O($\sqrt{\eta}$) of the longest path between two nearest neighbors in this optimal circuit, irrespective of the geographic spread of the $N$ parties. For example, in Fig. 3, the key rate scales quadratically as $d_{2,3}$ ($\eta\propto d$) rather than as the longest distance between two parties $d_{1,7}$.
2.2 Security aspects of multi-party quantum key agreement
When concerned with the security of quantum key agreement protocols, Eve, an eavesdropper, is the outside entity that can listen to all public communications. Eve may also attempt to manipulate the quantum channels and investigate correlations with the exchanged information. Theoretically, the Eve’s attacks are grouped into individual, collective, and coherent attacks [35]. Each protocol round can only be attacked by Eve separately in individual attacks. Her best action in this situation is to measure her quantum side information at each round, as it is presumed that she lacks quantum memory. We perform minimum error discrimination analysis [47] in Sec. 3, which can be considered an intercept-resend strategy for individual attacks. In collective attacks, Eve is thought to carry out the same attack for every protocol round, meaning that the quantum side information is independent and identically distributed (IID) across rounds. It is assumed here that Eve has a quantum memory, and as a result, she can execute a global operation at the end of the protocol and save her quantum side information for each round. We perform this analysis in Sec. 4 using source-replacement strategy [48, 49] to obtain the asymptotic secret key rate. Coherent attacks are the most general attack in which Eve’s capabilities are not limited, and they are generally used to study finite-size effects in the key.
3 Individual attacks and minimum error discrimination analysis
Initially we perform the simple minimum error discrimination analysis (on individual pulses) on the multi-party protocol to show its robustness compared to two-party protocol. Assuming that Eve controls the nodes AB, BC and the channels $l_{A}$, $l_{B}$, $l^{\prime}_{B}$ and $l_{C}$, she makes the announcements $\delta_{1}(\delta_{2})$ [47, 49]. Eve must effectively discriminate between the joint states that are correlated vs anti-correlated at node AB
@@ 3 Individual attacks and minimum error discrimination analysis
4 Collective attacks and the asymptotic key rate
To demonstrate the security against collective attacks and asymptotic key rate of the twin-field protocol, Ref. [49] introduces the source-replacement strategy [48], which is applied to Alice and Bob’s sources. Here we extend this to Charlie’s source. Then, we adapt the multi-party protocol to its equivalent entanglement-based protocol to verify the security against collective attacks in the asymptotic-key scenario. The key rate is then assessed to demonstrate the entanglement-based version’s security.
@@ 4 Collective attacks and the asymptotic key rate
5 ANSYS simulations and finite key rate
@@
Using simulations provides a convenient and powerful way to comprehend the workings of the QKD systems and the protocol implemented, when a full analytic treatment proves to be elusive. Only the key rate of the idealized protocol, which involves the quantum states based on attacks described in Sec. 2.2, is captured by our preceding theoretical analysis of security. The complete QKD security evaluation must consider not only the protocol scheme but also practical device flaws. Carefully evaluating the functional QKD systems and their defects requires taking into account the quantum states of the transmitted photons as well as the optical components. Simulations can be used to examine the theory-practice discrepancies and study countermeasures. The operational characteristics of the optical equipment, as well as environmental or Eve’s interference, are necessary to study a complete simulation. Initial QKD simulations used ideal assumptions and few optical components to numerically measure the performance [52, 53]. For the polarization-based QKD protocols such as BB84, extensive packages have been made available [54, 55, 56, 57]. Additionally, while polarization-based QKD methods are subjected to hacking attempts such as photon-number-splitting [58], hacking on measurement device-independent protocols like twin-field is different [59, 60, 61, 62]. We perform simulations of the protocol on the ANSYS Interconnect platform for finite key rate effects and describe the practical implementation with optical components.
The image in Fig. 4 shows a simulation for the multi-party twin-field protocol defined in Sec. 2.1. The rectangular selections show the modules of Alice, Bob and Charlie from top to bottom on the left side of Fig. 4.
@@
We consider 1 GHz pulse repetition rate with the line width of the laser as 5 kHz and analyze the key rate by performing multiple iterations of the simulations for varied distances $L$ in km with losses described by Eq. (29) ($L=l_{A}+l_{B}+l^{\prime}_{B}+l_{C}$ is fixed where $l_{A}=l_{B}=l_{B}^{\prime}=l_{C}$ in FIBER_1, FIBER_2, FIBER_3, FIBER_4 of Fig. 4). We obtain the secret key rate in the log scale, which closely matches the behavior of the result from the previous section, as shown in Fig. 10. The final secret key rate (skr) is obtained from the sifted key rate (kr) based on the Holevo bound from Eq. (19), where $\hbox{skr}=\hbox{kr}\times(1-\chi)$. For a finite number of pulses sent ($N=10^{9}$), the key rates are lower than the asymptotic key rate from Eq. (26), as expected. Further, in Fig. 10, we observe that by varying the mean photon number, $\mu=0.2$ shows optimal secret key rate compared to $\mu=0.5$ (considering $\mu_{1}=\mu_{2}=\mu$ in Eq. (26)). The secret key rate was more than $10^{-6}$ per pulse (1 kbps) for $L=$ 250kms.
Figure 10: Secret key rate from simulations compared with asymptotic key rate from Eq. (26). The optimal value of mean photon number $\mu$ for higher secret key rate is around 0.2 with finite pulses.
@@ 6 Difficulty in practical implementation
Practically putting in place a twin-field protocol between multiple parties presents technological difficulties since active phase stabilization is necessary to interfere with coherent pulses due to random phase fluctuations [10, 44, 45, 46]. It is worth noting that this problem presents the main bottleneck in realizing twin-field cryptography, and calls for unique technology developed to address it. Here we mention two such methods: (a) in one case, the wavelengths of two users’ operating two separate lasers are locked to estimate and compensate for phase drift [44, 46]; (b) In another case, phase fluctuation was overcome by matching the modes of two independent lasers using the laser injection technique and the phase post-compensation approach [45]. The central nodes AB and BC function as gigantic Mach-Zehnder interferometers for interference measurement, as shown in Fig. 2. Stabilizing an unbalanced Mach-Zehnder interferometer with widely different path lengths is challenging. So, the setup would necessitate that users Alice-Bob (Bob-Charlie) have an identical geographic distance to the central node AB(BC) for the interference of pulses ideally, whereas, in the real world, there could be significant variations in the geographic distances between various parties and the nodes. A more practical method to accomplish this is via sending-or-not-sending, a unique type of twin-field protocol where data is encoded based on the decision to send in a coherent state (sending) or a vacuum state (not sending), and it has tolerance to phase misalignment errors [65].
@@ 7 Conclusion
We propose and examine a technique that expands on the two-party twin-field quantum key distribution to obtain a protocol for the concurrent multi-party quantum key agreement. For the proposed protocol, we carry out the asymptotic key rate analysis from a security proof based on the technique of entanglement-based source replacement. We simulated the protocol on the ANSYS Interconnect platform by assessing its practical implementation. These simulations can be augmented by including realistic experimental issues such as phase misalignment errors and the use of decoy states with varying intensities. This exercise can be extended to extensively study the variant of the present protocol based on the sending-or-not-sending twin-field protocol [65]. These simulations can also be used to quantify the information lost by the quantum hacking efforts [3, 5, 61]. Trojan-horse attack [67, 68, 69, 70, 71] and backflash attack [72, 73] were studied on distributed-phase-reference QKD simulations using ANSYS, which might also be relevant to twin-field implementation [74].